OpenStudy (anonymous):

what is 5a-b - 4a-2b in simplest form ----- ------ a^2-b^2 a^2-b^2

OpenStudy (anonymous):

hi :) so what happens is we look at the denominator. They are the same, which is a great sign. So, we simplify the top: 5a - b - (4a - 2b) = a - b + 2b = a + b Take note that - (4a - 2b) is actually -4a + 2b! Then, we notice that the denominator can be simplified into the form (a+b)(a-b). Which is a great thing, since we can then cancel them out. And so, we are left with 1 over (a-b). In other words, 1 ------ a-b have a nice day! :)